Dakota Inc. and Jersey & Company are two large companies that manufacture and sell equipment used...

  1. Dakota Inc. and Jersey & Company are two large companies that manufacture and sell equipment used in the construction, mining, agricultural, and forestry industries. The companies reported the following data (in millions) for two recent years:

    Dakota Jersey
    Year 2 Year 1 Year 2 Year 1
    Net income $2,172 $3,705 $1,930 $3,177
    Average number of common shares outstanding 594 599 334 363

    a. Determine the earnings per share in Year 2 and Year 1 for each company. Round your answers to two decimal places.

    Year 2 Year 1
    Dakota $ per share $ per share
    Jersey $ per share $ per share
Answer #1
Year 2 Year 1
DAKOTA (2172/594) = 3.66 (3705/599) = 6.19
JERSEY (1930/334) = 5.78 (3177/363) = 8.75

Note :

I have rounded some numbers. Comment if you face any issues